Flying Cloud Co. has the following operating data for its manufacturing operations: Unit selling price $210 Unit variable cost $115 Total fixed costs $783,000 The company has decided to increase the wages of hourly workers which will increase the unit variable cost by 10%. Increases in the salaries of factory supervisors and property taxes for the factory will increase fixed costs by 4%. If sales prices are held constant, the next break-even point for Flying Cloud Co. will be a. increased by 1,208 units b. decreased by 1,510 units c. increased by 1,812 units d. increased by 1,510 units
Answer #1

SOLUTION

Correct answer is increased by 1,510 units thus, Option D is correct.

Existing  break-even point for Flying Cloud Co. = Fixed Cost / Contribution Margin Per Unit

= Fixed Cost / (Sales Price Per Unit - Variable Cost per Unit)

= $783,000 / ($210 - $115)

= $783,000 / $95

= 8,242 units

Revised Variable cost = $115 * 110% = $126.50

Revised Fixed cost = $783,000 * 104% = $814,320

Hence, Revised break-even point for Flying Cloud Co. = Fixed Cost / Contribution Margin Per Unit

= Fixed Cost / (Sales Price Per Unit - Variable Cost per Unit)

= $814,320 / ($210 - $126.50)

= $814,320 / 83.50 = 9,752 units

Hence, Increase = 9,752 - 8,242 = 1,510 units